You plan to buy a house in December 2019. The sale price is $385,000. You need to pay 20% down payments and borrow an additional 80% from Bank of America with a 25-year, 3.0% fixed-rate mortgage loan. You are expected to pay an equal MONTHLY payment starting from January 2020 for a total of 25 years.

(1) Calculate your expected monthly mortgage payment.

(2) Develop the 2020~2044 amortization table for your mortgage loan.

(1)   When you prepare your 2020 tax filing, what is the total mortgage interest payment that you can consider for the tax deduction?
Answer #1

1.

1460.57

2.

308000 25 2 Loan principal 3 Term(years) 4 Term (months) 5 Annual interest 6 Monthly interest 7 Monthly payment 8 If Quarterl

=385000*(1-20%) =B3*B8 0.03 =B5/B8 =-PMT(B6,B4,B2) 12 2 Loan principal 3 Term(years) 4 Term (months) 5 Annual interest 6 Mont

3.

9125.10